OpenStudy (anonymous):

Medal and fan reward! help appreciated! evaluate -3a divided by B for a=5/6 b=1/6 Please show your substitutions.

OpenStudy (anonymous):

-3a divided by b a=5/6 b=1/6

OpenStudy (anonymous):

-3a/b = -3*(5/6)/(1/6) = -(3*5)*[(1/6)/(1/6)] = -3*5*[1] = -15

OpenStudy (anonymous):

r u 100 percent sure?